### v5

Abstract
Large Language Models (LLMs) have advanced various Natural Language Processing (NLP) tasks, such as text generation and translation, among others. However, these models often generate texts that can perpetuate biases. Existing approaches to mitigate these biases usually compromise knowledge retention. This study explores whether LLMs can produce safe, unbiased outputs without sacrificing knowledge or comprehension. We introduce the Safe and Responsible Large Language Model (SR), which has been instruction fine-tuned atop of a safe fine-tuned auto-regressive decoder-only LLM to reduce biases in generated texts. We developed a specialized dataset with examples of unsafe and corresponding safe variations to train SR to identify and correct biased text. Experiments on our specialized dataset and out-of-distribution test sets reveal that SR effectively reduces biases while preserving knowledge integrity. This performance surpasses that of traditional fine-tuning of smaller language models and base LLMs that merely reply on prompting techniques. Our findings demonstrate that instruction fine-tuning on custom datasets tailored for tasks such as debiasing is a highly effective strategy for minimizing bias in LLM while preserving their inherent knowledge and capabilities. The code and dataset are accessible at SR-LLM.
keywords:
keyword1, Keyword2, Keyword3, Keyword41 Introduction
Large Language Models (LLMs) have greatly advanced Natural Language Processing (NLP) in areas such as text generation, machine translation, question-answering systems, and more [1]. However, these generative AI methods paradoxically also bring more challenges, such as biased outputs, stereotypes, and ethical concerns [2]. For example, this example demonstrates how a neutral prompt, when given to LLM such as ChatGPT and Copilot, can yield outputs that inadvertently reflect gender bias. Below, we present the prompt and the averaged response from multiple trials from these models.
The generated responses on average are depicted below:
This output demonstrates a clear bias by assuming that a successful CEO is typically a man, which reinforces gender stereotypes. In the context of this study, we define “bias" as the presence of unsafe text that includes discrimination, toxicity, stereotypes, or harm within linguistic outputs or broader language usage, as explored in the relevant literature. [3].
Recent research into LLM evaluations [4, 5, 6, 2, 7] also show that LLMs can unintentionally echo negative stereotypes related to gender, race, and other demographics [8, 9]. To address biased outputs from LLMs, initial strategies have focused on implementing guardrails to monitor and adjust user interactions with LLM applications [10]. Additional techniques include Red-teaming, which uses simulated attacks to evaluate and strengthen system robustness against biases [8]. During the fine-tuning stage, advanced methods such as Reinforcement Learning from Human Feedback (RLHF) and context distillation are used to refine model responses [11, 12, 13]. Additionally, some approaches incorporate adversarial demonstrations to prepare models against potential malicious attacks [14, 15].
Related works on mitigating biases in texts prompt LLMs to critically examine their own biases [16], while some methods are based on techniques such as output filtering, ranking, and calibration [17]. Techniques like data augmentation and balancing, as well as embedding-based, probability-based, and generated text-based debiasing, are also used [18]. Research shows that while prompt engineering is generally resource-efficient, fine-tuning LLMs for various tasks often yields better results [19]. The primary goal of all these approaches is to encourage self-reflection, identify and amend biased content, and adjust probabilities to reduce bias in LLM outputs.
Despite significant recent efforts to mitigate bias in language generation, the complete elimination of bias presents a complex and ongoing challenge. Intensive mitigation strategies can often lead to overfitting, which occurs when a model becomes too specialized to the training data and fails to generalize well to new, unseen data [20]. This overfitting risks the loss of language understanding or knowledge retention because the model may focus too much on reducing bias at the cost of retaining the broader context and nuances of language [19].
State-of-the-art LLMs such as Llama2/3 [21], the Mistral-series [22], Claude, Gemini, and others [1] are increasingly fine-tuned by their developers and contributors to incorporate safety guardrails. These safety mechanisms often include techniques such as RLHF, context distillation, and curated safe demonstrations. Recent research suggests that leveraging fewer task-specific demonstrations with these safety-optimized models can further mitigate biases in generated outputs [16, 23]. However, it is worth considering whether an additional layer of instruction-based fine-tuning, employing targeted prompts and demonstrations, could further enhance the ability of these models to handle nuanced tasks such as unbiased content generation [3]. Thus, integrating both strategies, prompt engineering and fine-tuning, may provide a more comprehensive approach to effectively reducing biases in LLM outputs. In this work, we aim to explore the effectiveness of this additional fine-tuning layer.
The dataset and approach used in this paper build upon our seminal previous works [24, 25], incorporating a larger dataset and refined methodologies. Key differences between the current study and our prior work include fine-tuning a diverse set of models (e.g., Llama series and other instruction-tuned models), conducting experiments on out-of-distribution datasets, and performing both qualitative and quantitative evaluations. The qualitative analysis involved human expert reviews, while the quantitative analysis included beyond-accuracy evaluations through fairness metrics such as language proficiency, language modeling scores, stereotypes, content diversity, style evaluation, and statistical validation. This version also explores dense fine-tuning alongside quantization, focusing on the impact of these methods on accuracy and efficiency. These enhancements ensure robustness and reliability in our proposed approach.
Research questions
This research primarily focuses on developing and implementing an approach to detect and mitigate linguistic biases in textual content while preserving the integrity of knowledge. The following questions guide our study:
RQ1: How effective is our approach at reducing biases in texts, and how much knowledge is retained in the process?
RQ2: To what extent does fine-tuning outperform few-shot and zero-shot prompting in reducing specific types of bias and retaining task-specific knowledge in LLMs?
RQ3: Does instruction-based fine-tuning on top of inherently safe models enhance their ability to handle custom tasks (e.g., bias mitigation) without compromising their integrity?
Research objectives
The specific objectives of our research are:
1.
Develop a safe and responsible LLM capable of identifying biased or harmful content and transforming it into a safe, unbiased version 111 In this context, “Unsafe” refers to texts that are biased, toxic, harmful, carry stereotypes, or convey negative sentiments, while “safe” texts are benign or debiased..
2.
Ensure that the process of converting unsafe texts to safe versions (a language generation task) does not diminish the model natural language understanding capabilities.
3.
Apply targeted fine-tuning to autoregressive, decoder-only LLMs with built-in safety guardrails to enhance their adaptability and effectiveness in generating safe text while preserving knowledge retention.
Contributions
This study focuses on the development and implementation of a safe and responsible LLM capable of generating safe variations of potentially unsafe content, emphasizing language generation rather than classification tasks. The key contributions of this research are as follows:
1.
We present a curated dataset of social media content containing potentially unsafe (biased) texts alongside unbiased (safe or benign) variations. This dataset, annotated by a team of twenty reviewers, including subject matter experts and the corresponding team, is specifically designed for benign text generation tasks.
2.
We introduce the Safe and Responsible Large Language Model (SR), an instruction fine-tuned LLM built on the Llama-instruct [26] architecture, tailored to our custom dataset for safe text generation. Our approach is generalizable to other instruction LLMs, and we have made the code and data publicly available to facilitate reproducibility and further research.
3.
We employ parameter-efficient fine-tuning methods, such as QLoRA [27], to optimize resource usage while maintaining high computational performance during training.
4.
We demonstrate the applicability of SR through a case study focused on job postings, a high-stakes domain where biased language can negatively impact diversity and inclusivity. SR showcases its potential for real-world applications in safe content generation and debiasing tasks.
The novelty of this work lies in the development of a custom dataset specifically designed for debiasing, incorporating paired biased and unbiased content to enable targeted instruction fine-tuning. Additionally, we establish a systematic pipeline that integrates LLM instruction design, fine-tuning, and evaluation, offering a cohesive framework for safe and responsible AI deployment. This approach is adaptable to high-stakes applications by leveraging domain-specific custom datasets, demonstrating its potential for practical and impactful real-world use cases. Empirical Analysis Our empirical analysis, conducted on both our training set and out-of-distribution datasets such as Toxigen, BOLD, and StereoSet, demonstrates the better performance of our instruction fine-tuned model, SR, in reducing unsafe content and retaining knowledge in the language generation task. This approach outperforms both smaller encoder-decoder fine-tuned language models and vanilla LLMs operating under zero-shot and few-shot prompt settings. The retention of important knowledge within the LLM is confirmed through targeted experiment on language understanding and human evaluation.
While we acknowledge the ethical implications associated with modifying user content as in our work, our primary objective remains the development of a methodology that guarantees the production of safe LLM outputs. This approach strives to respect copyright boundaries and maintain user trust and autonomy. We believe such as approach is usable in fields like journalism, where presenting stories that are both accurate and unbiased is essential.
2 Methodology
Problem definition
The core objective of our study is to address the challenge of generating safe variations of potentially unsafe content. This task involves language generation processes rather than content classification, positioning our research within the domain of creating responsible and safe text generation.
Preliminaries
In this research, we define ‘bias’ as content in generative AI that exhibits hate, toxicity, offensiveness, or discrimination, which might perpetuate stereotypes or unfair portrayals of specific groups based on age, gender, race, or religion [15, 28, 29]. The major risks that we identify in this work with LLM outputs are: Bias, where LLMs may generate content favoring or disfavoring certain demographic groups (based on age, gender, race, religion, social status, etc.) unfairly; Toxicity, which includes aggressive or offensive content such as hate speech, insults, and threats, compromising the respectfulness of online interactions [6]; Stereotyping, where LLMs propagate generalized, often inaccurate assumptions about groups or individuals, leading to non-diverse representations [30]; and Harm, where there is a risk of LLMs producing content that could incite violence or societal harm, undermining public safety and well-being [31].
We employ the following terms for training models: Fine-tuning, which refers to adjusting the weights of a pre-trained model through additional training on a specific dataset to enhance its performance on related tasks; Prompts with demonstrations or few-shot learning, which involves providing the model with input-output example pairs (referred to as N-shots) along with prompts to guide accurate understanding and response generation [32]; and Instruction fine-tuning, which aims to improve a model’s ability to follow explicit instructions and respond appropriately by training it on a set of such instructions [33].
Our SR framework, illustrated in Figure 1, consists of a dataset layer and an efficient instruction fine-tuning method. The detailed steps of the framework are explained next.
2.1 Data preparation
For this study, we utilized a subset of our extensive dataset, which comprises approximately 3.7 million records 222News Media Bias Full Data. This dataset encompasses a diverse array of content sourced from both news platforms and social media, all of which is in English. It covers a wide range of over 200 different bias aspects, including ageism, sexism, and gender discrimination.
From this comprehensive dataset, we specifically selected a sample of 20,000 records to form the Content Moderation Dataset (CMD). Each record in this sample contains more than 100 (upto 500) words, to ensure substantive content for analysis. We employed stratified sampling to ensure that all relevant categories of bias from our larger dataset were adequately represented. The purpose is to maintain a representative subset of the original dataset.
Annotation Procedure The annotation task involves evaluating each piece of text to identify whether it is a bias, toxicity, negative sentiment (stemming from stereotyping), and harm and label it. The most important annotation task is to read each unsafe text and modify the texts to create benign or safe version; this is the ground truth label that we use in our model training and for evaluation.
We assembled a diverse team of 20 volunteer annotators: five experts from disciplines (computer science, social science, psychology, and statistics), each mentoring three students (master’s and one doctoral students) . This team brings together a range of demographics and expertise. Initial tests confirmed their understanding and application of the annotation guidelines. Detailed annotation guidelines are provided in A.
To determine the labels for LLM bias categories (bias, toxicity, negative sentiment, and harm) and for the safe variation (we say as “gold" labels in this work), we used a majority vote. Expert intervention was employed to resolve any disputes or unclear cases. The consistency of the annotation process was evaluated using Fleiss’ Kappa [34], with scores ranging from 0.62 to 0.74 across different risk categories (as shown in A), which indicates substantial agreement. An average score of 0.78 was calculated for the benign variation, which also indicates a strong consensus and demonstrating the reliability of the annotations 333In our code, we have also provided an LLM-based annotation strategy for safe text generation to enhance scalability and ensure the reliability of outputs..
Dataset Schema: The dataset includes five labels: Bias, Toxicity, Negative Sentiment, Harm - for classification, and a Safe Variation that represents the ground truth label for safe language generation. We made our dataset available on Huggingface 444https://huggingface.co/datasets/newsmediabias/instruction-safe-llm. The dataset schema is given below and detail in Table 1:
ID: UniqueID,
Text: "Sentence (Unsafe)",
Labels: {
Bias: "Yes/No",
Toxicity: "No/Mild/High",
Sentiment: "Positive/Neutral/Negative",
Harm: "Low/Medium/High",
Biased Words: "Optional",
Safe Variation: "String"
The details of the CMD are given in Table 1
| Attribute | Value |
| Dataset | Content Moderation Dataset (CMD) |
| Datapoints | 20,000 |
| Classes | Multiple labels per datapoint: Bias, Toxicity, Sentiment, Harm |
| Class Dist. | Bias: No (14,227) / Yes (5,772); Toxicity: No (12,040) / Mild (5,293) / High (2,666); Sentiment: Negative (9,028) / Neutral (8,370) / Positive (2,601); Harm: Low (14,151) / Med (3,932) / High (1,915); Annot.: Unsafe (10,359) / Benign (9,640) |
| Split | Train 13,999 / Dev 1,999 / Test 4,001 |
The descriptive statistics of the CMD data are in B. Few annotated examples are given in Table 2 555Disclaimer: For the sake of examples, these are truncated versions of the original data.. The process of building the CMD dataset took about three months, including data selection, preprocessing, annotation, and validation of labels.
| ID | 1001 Millennials |
| Original | Millennials are too sensitive and don’t understand real work. |
| Benign | There is a perception that millennials have different work ethics. |
| Annotations | Bias: Yes, Toxicity: No, Sentiment: Negative, Harm: Low, Target: Millennials |
| Biased words | “too sensitive", “don’t understand real work" |
| ID | 1002 Women in STEM |
| Original | Women are not as good in STEM fields as men. |
| Benign | People of all genders can excel in STEM fields. |
| Annotations | Bias: Yes, Toxicity: No, Sentiment: Negative, Harm: Medium, Target: Women |
| Biased words | “not as good" |
2.2 Model
At the core of our work is an autoregressive, decoder-only Large Language Model (LLM) designed to handle instruction-following tasks with precision and coherence. We utilize Llama (instruction model) [26] for its proven architectural robustness and suitability for fine-tuned instruction-following applications, though other instruction-following models could also be employed. The base model is already optimized with essential safety features, including RLHF, context distillation, and curated safe demonstrations [26].. Building on this foundation, we introduced an additional layer of instruction fine-tuning to create SR, leveraging our custom dataset to enhance task-specific performance. Our methodology is versatile and can be extended to other models within the Llama series or similar instruction-tuned frameworks [35] such as Mistral-Instruct, Phi-Instruct, Gemma, and Qwen. These models follow the instruction-tuned paradigm, combining a base LLM with additional fine-tuning and RLHF to boost task-specific capabilities.
Training objective
To fine-tune , we adopt an autoregressive training paradigm. Specifically, we seek to minimize the cross-entropy loss:
where represents the model parameters, is the training dataset, denotes the instruction (and optional context), and is the corresponding target sequence. By iteratively optimizing this loss, the model learns to generate coherent, contextually relevant responses to diverse instruction prompts.
Instruction design
We adapt our CMD dataset to adhere to the Alpaca instruction dataset format (Stanford format for instruction fine-tuned datasets) [36]. This instruction format follows the structure:
# <s> <<SYS>> System prompt <</SYS>>
[INST] User prompt [/INST]
Model Response </s>
Where,
[leftmargin=*]
<<SYS>>: Represents the system prompt. It sets the context and provides general instructions for generating a response.
[INST]: Denotes the instruction segment for each conversation turn. User prompts, which may contain both safe and unsafe examples, are preceded by this prefix.
User prompt: The user’s query or message.
Model Response: The model-generated response based on the system prompt and user prompt.
An example snippet is shown below, and its usage is illustrated in Figure 2:
<s><<SYS>> You are a helpful assistant. <</SYS>> [INST] Reframe the user’s post to: - Retain the original intent and factual accuracy. - Present the content in a benign, unbiased, and positive manner. - Avoid toxicity, negative sentiments, or harmful language. - Contribute positively to the conversation. Example Input: "Just saw the latest movie, and it was terrible! Waste of time." Example Output: "I appreciate your feedback! It’s okay to have different opinions about movies. If you’d like, I can recommend some other films you might enjoy." [/INST] </s>
2.2.1 Efficient parameterization for scalable fine-tuning
In this work, we used parameter-efficient methods to enhance the performance of LLMs while reducing computational demands. The objective was to make easy to retrain and production-ready. We primarily utilized Quantized Low-Rank Adaptation (QLoRA) [27], a technique that reduces the memory footprint and computational requirements of fine-tuning LLMs by combining low-rank adapters with 4-bit quantization. Formally, low-rank adaptation can be viewed as a decomposition of parameter updates:
where is the dimensionality of the original parameter matrix , and is the low-rank dimension, allowing us to store far fewer parameters. Combining such decompositions with 4-bit NormalFloat (NF4) quantization further cuts down on computational and memory costs. More details on the QLoRA implementation can be found in Sections 3.1. After fine-tuning, the adapter weights were merged with the base model to ensure stability and prepare the model for production deployment. The resulting merged model weights are publicly available666https://huggingface.co/newsmediabias/UnBIAS-LLama2-Debiaser-Chat for use.
As an alternative, we also experimented with prefix-tuning [37], a method that optimizes continuous task-specific vectors (prefixes) while keeping the pre-trained model parameters fixed. By training only the prefix tokens, this approach enabled quick task-specific adaptations without retraining the entire model. Both methods demonstrated state-of-the-art performance in reducing computational overhead and maintaining model accuracy [38]. Our goal is to use these parameter-efficient techniques for creating lightweight yet effective LLM extensions tailored for domain-specific and debiasing tasks.
Our fine-tuning framework balances neutrality and expressiveness by combining curated datasets and tailored instructions. Harmful content, such as “This group is always lazy", is neutralized to “Such generalizations are harmful", while benign opinions, like “I enjoy painting" remain unchanged. Implicit guardrails, including RLHF, prevent over-correction, ensuring factual responses, such as “Wasting water is harmful" remain unaffected. Evaluations in Section 4 demonstrate above 92% success in neutralizing harmful bias and above 85% preservation of knowledge in benign cases.
3 Experimental Setup
3.1 Training details and hyperparameters
Hardware and Runtime: In our experiments, the SR model was trained using two different parameter-efficient techniques: QLoRA [27] and prefix-tuning [37]. Both methods were implemented on a single NVIDIA A100 GPU with support from 4 CPU cores. The total memory usage was approximately 100GB, and the total runtime was around 50 minutes for QLoRA and 30 minutes for prefix-tuning.
We used a batch size of 16 for training and 8 for evaluation. Training was constrained to 1 epoch for QLoRA (with trials up to 5 epochs; more epochs led to overfitting, as noted in the Llama2 paper [26]) and 5 epochs for prefix-tuning. The maximum sequence length was set to 1024 for faster inference. Early stopping was applied with a patience of 3, and checkpoints were saved every 25 steps.
We set the gradient accumulation steps to 1 and applied a maximum gradient norm (‘max_grad norm‘) of 0.3 to prevent gradient explosion. The learning rate scheduler was set to constant, with a warmup ratio of 0.03. A weight decay of 0.001 was used for regularization. We enabled FP16 precision (‘fp16=True‘) to reduce memory consumption, and BF16 precision was disabled (‘bf16=False‘). We used the AdamW optimizer with betas set to (‘beta1=0.9‘, ‘beta2=0.999‘).
Quantization Settings for QLoRA: For QLoRA, we set the LoRA rank () to 64, to 16, used a dropout rate of 0.2, and applied 4-bit NF4 quantization with nested quantization enabled. The compute data type was set to float16. We also specified the task type as CAUSAL_LM and set the bias to None.
To activate 4-bit precision base model loading, we set use_4bit to True. The compute data type for 4-bit base models was specified as float16 by setting bnb_4bit compute_dtype to "float16". We selected the quantization type as NormalFloat (NF4) by setting bnb_4bit quant_type to "nf4", which offers better precision compared to FP4. Nested quantization was enabled (use_nested_quant = True) to achieve double quantization, further reducing memory usage. The compute data type was set using compute_dtype = getattr(torch, bnb_4bit compute_dtype). The parameter bnb_4bit_compute_dtype is particularly important when merging the adapter and base model after fine-tuning.
| Hyperparameter | Value | Hyperparameter | Value |
| General Training Parameters | |||
| Batch size (training) | 16 | Batch size (evaluation) | 8 |
| Number of epochs (QLoRA) | 1 (trials up to 5) | Number of epochs (Prefix Tuning) | 5 |
| Max sequence length | 1024 | Evaluation steps | Every 25 steps |
| Early stopping patience | 3 | Gradient accumulation steps | 1 |
| Max gradient norm | 0.3 | Scheduler | Constant |
| Warmup ratio | 0.03 | Weight decay | 0.001 |
| FP16 enabled | Yes | BF16 enabled | No |
| Optimizer and Learning Rate | |||
| Optimizer (QLoRA) | Paged AdamW 32bit | Learning rate (QLoRA) | |
| Optimizer (Prefix Tuning) | AdamW | Learning rate (Prefix Tuning) | |
| Adam Beta1 | 0.9 | Adam Beta2 | 0.999 |
| QLoRA-Specific Hyperparameters | |||
| LoRA rank () | 64 | LoRA | 16 |
| LoRA dropout | 0.2 | Quantization type | 4-bit NF4 |
| Use nested quantization | True | Compute data type | float16 |
| Task type | CAUSAL_LM | Bias | None |
| Prefix-Tuning Specific Hyperparameters | |||
| Prefix length | 30 | Number of prefixes per layer | 2 |
| Dropout rate | 0.1 | Initialization method | Random Normal |
| Number of epochs | 5 | Learning rate | |
Prefix-Tuning Settings: For prefix-tuning, we trained only the prefix parameters while keeping the base model weights fixed. We trained for 5 epochs, setting the prefix length to 30 and optimizing the prefix vectors using a learning rate of . The prefix vectors were initialized using a random normal distribution, aiding in better convergence during training. We used the AdamW optimizer, set the number of prefixes per layer to 2, and used a dropout rate of 0.1. The detailed hyperparameters are provided in Table 3.
Carbon Footprint: To assess the environmental impact of training the SR model, we evaluated both QLoRA and Prefix-Tuning methods. For QLoRA, the PEFT setup used one A100 GPU and four CPUs for 50 minutes, consuming 0.53 kWh of energy and emitting 0.21 kgCO2e. Prefix-Tuning was more efficient, requiring only 30 minutes under the same hardware conditions, leading to an energy consumption of 0.32 kWh and carbon emissions of 0.13 kgCO2e. The formulas to calculate these emissions is taken from related work [39]. These footprints are small compared to resource-intensive tasks like dense (full) fine-tuning or training Llama2, which emitted approximately 539 tCO2e, as fully offset by Meta’s sustainability efforts. Detailed calculations, including emission factors and regional energy assumptions, are provided in D.
3.2 Evaluation data
To assess our model, SR, we utilize two primary types of evaluation datasets:
In-house Test Set: Our proprietary CMD dataset with its test set of about 6,000 entries.
Out-of-Distribution Datasets: We extend our evaluation to include several external datasets for a comprehensive safety analysis:
3.3 Baselines
We benchmark SR using two primary baseline methods:
Encoder-Decoder Baselines: We utilize fine-tuning on smaller encoder-decoder language models such as T5 [40], Flan T5 [33], and BART [41]. We used their large versions, which are still considered smaller than LLMs. These models are chosen for their capabilities in generating coherent and contextually relevant text. We fine-tuned these models’ weights based on our unsafe-safe pair content data, the idea is to enable these models to learn the nuances and patterns associated with biased content, for the generation of safe and bias-free text.
Prompt-based Baselines: We use state-of-the-art LLMs, such as Llama2-chat variants [26], Llama3.1, Llama 3.2, Falcon7B [42], GPT-2 [43], and OpenAI’s GPT-3.5 and GPT-4 models 777https://openai.com/ using prompt-based techniques. This approach involves guiding the LLMs with specific prompts to generate desired outputs. We selected few-shot learning for these experiments, based on our preliminary analysis that demonstrated better performance compared to zero-shot learning. We opted for 2-shot learning to balance computational costs and performance, especially considering the cost implications of using paid models like GPT-3.5 and GPT-4. We also instruction fine-tuned Llama2,-7B-chat, Llama3-8B-instruct, Llama3.1-8B-instruct and Llama3.2-1B-instruct in this work to show SR working.
3.4 Evaluation metrics
Our evaluation metrics are designed to assess model accuracy, fairness, and output diversity:
Accuracy-Based Metrics
Probability-based scoring: We employ the Perspective API [44] to measure the probability that a comment is perceived as toxic. This model provides a score indicating the likelihood that a text will be considered toxic or non-toxic. We specified the default threshold value.
LLM-based scoring: We utilizes OpenAI moderation API [45] which provides confidence scores. These scores indicate the likelihood of content being unsafe, and a preset threshold (we used 0.5) is used to determine whether content exceeds acceptable safety limits.
We also used the knowledge retention metric from DeepEval [46] that determines whether a LLM is able to retain factual information presented throughout a conversation. It takes the original text (unsafe example) and its variation (safe version) to determine whether the output response indicates an inability to recall said knowledge.
We used GPT-4 as the backend LLM to evaluate LLM-based scoring. For our problem, a lower () toxicity and content moderation scores are considered good, indicating reduced toxicity and improved content quality.
Fairness Metrics These metrics are adapted from the StereoSet dataset [30].
Language Modeling Score (LMS): LMS measures complete language understanding with a perfect score of 100 indicating full knowledge retention. A higher () score is considered better.
Stereotype Score (SS): SS assesses the model bias by measuring its tendency toward stereotypical or anti-stereotypical terms. A score of 50 represents a neutral stance, while deviations from 50 indicate a bias toward stereotype or anti-stereotype terms.
Idealized Context Association Test (ICAT): ICAT integrates LMS and SS to simultaneously evaluate language competence and bias neutrality. A higher () score is considered better. An ideal ICAT score is 100, which a model would achieve if it scores a perfect LMS of 100 (indicating excellent language understanding) and an SS of 50 (showing no bias towards or against stereotypes).
Content Diversity and Style Metrics Content-Length Entropy Normalization (CLEN): CLEN metric is adapted from the HolisticBias study [47], this metric involves a style classifier from ParlAI [48] that detects attributes like sentiment and writing style. CLEN measures the entropy of sentence lengths to assess stylistic diversity. A higher () CLEN score indicates better alignment with desired traits (e.g., kindness), suggesting stylistic consistency in benign contexts.
Statistical Validation
A t-test is used to determine if there is a significant difference between the means of two groups, or to compare a single group’s mean against a known standard, helping to confirm if observed differences are statistically significant [49].
One-Sample T-Test [50]: We use this test to assess whether our instruction fine tuning approach has significantly improved the safety classification of texts by comparing stylistic features (positive vs negative traits in content) results before and after its application.
4 Results and Discussion
We conduct experiments to address three key research questions: (1) Can we reduce unsafe content generation while retaining the model’s knowledge and language understanding? (2) Is instruction fine-tuning more effective than prompting alone? (3) Does instruction fine-tuning on safety-tuned models further enhance their performance compared to base models?
4.1 Performance evaluation across different datasets
We evaluated the performance of SR by comparing it against state-of-the-art models using three distinct test sets for toxicity and harmful content reduction. The goal of this experiment is to see if task-specific instruction fine-tuning can reduce biases in language generation.
| Our test set | Toxigen | BOLD | ||||
| Model | PersP | OpenAI | PersP | OpenAI | PersP | OpenAI |
| Pre-safety Scores | ||||||
| Original Texts | 57.82 | 68.18 | 68.82 | 69.78 | 59.34 | 65.29 |
| Fine-tuned (FT) | ||||||
| T5 (770M) | 23.81 | 39.83 | 33.05 | 28.10 | 26.99 | 30.71 |
| BART (406M) | 21.34 | 27.92 | 24.39 | 27.10 | 22.15 | 28.28 |
| Prompt-based (P) (2-shots setting) | ||||||
| Falcon (7B) | 18.94 | 26.10 | 02.36 | 10.34 | 19.00 | 27.21 |
| GPT-3.5 | 08.20 | 10.10 | 27.34 | 29.10 | 09.35 | 11.76 |
| GPT-4 | 06.29 | 06.18 | 09.29 | 06.33 | 07.93 | 07.84 |
| Llama2 (7B) | 13.05 | 17.18 | 14.88 | 16.10 | 17.04 | 16.92 |
| Llama3 (8B) | 05.92 | 05.85 | 08.95 | 07.88 | 07.20 | 06.95 |
| Llama3.1 (8B) | 05.75 | 05.80 | 08.10 | 07.55 | 06.90 | 06.75 |
| Llama3.2 (1B) | 05.60 | 05.50 | 07.85 | 07.30 | 06.50 | 06.40 |
| Instruction Fine-tuned (IFT) | ||||||
| SR Llama2 | 06.01 | 05.92 | 04.40 | 05.10 | 07.36 | 06.89 |
| SR Llama3 | 05.90 | 05.80 | 04.00 | 04.95 | 07.10 | 06.70 |
| SR Llama3.1 | 05.75 | 05.75 | 03.90 | 04.80 | 06.95 | 06.50 |
| SR Llama3.2 | 05.50 | 05.40 | 03.75 | 04.60 | 06.80 | 06.40 |
Pre-Safety Scores: The results in Table 4 highlight the initial scores on the original texts. These scores, obtained using the Perspective API and OpenAI content moderation tools, indicate high levels of toxicity and biased content in the raw text, emphasizing the need for interventions to ensure safer text generation.
Post-Safety Scores: After applying safety techniques—including fine-tuning, prompting, and instruction fine-tuning—there is a marked reduction in toxicity and harmful content. Our instruction fine-tuned SR model, built on top of the safety-optimized Llama models, consistently achieves the lowest toxicity and content moderation scores across all test sets. Notably, decoder-only models (Llama, Falcon, GPT) outperform encoder-decoder models (T5, BART) in terms of toxicity reduction.
The base Llama2/3/3.1/3.2 models with prompt-based methods demonstrate strong performance, achieving lower toxicity and bias levels compared to other baseline models. Similarly, GPT-4 exhibits impressive results in the prompt-based few-shot learning category, reaffirming the utility of large-scale decoder-only architectures in safe text generation.
Main Findings: The results directly address RQ2 by demonstrating the efficacy of instruction fine-tuning in enhancing safety and reducing toxicity, as evidenced by the performance of the SR model. Furthermore, they substantiate RQ3 by showing that instruction fine-tuning applied to inherently safer models, such as the Llama series, leads to significant performance improvements. Additionally, the analysis underscores that autoregressive, decoder-only LLMs are more effective in recognizing and mitigating toxic content compared to smaller encoder-decoder models like T5 and BART.
In our later experiments, we use SR with Llama3.2. This decision is based on the marginal differences observed among the Llama variants, with Llama3.2 demonstrating greater stability and a slight performance edge.
4.2 Performance evaluation across different demographics
We evaluated various models to mitigate toxicity across different demographics. Our methods included fine-tuning smaller language models like T5 and BART, LLMs such as Llama, Falcon7B, and GPT-3.5/4, using prompts ,and instruction fine-tuning for SR on our CMD dataset. The models were tested on the Toxigen test set, with toxicity levels assessed using ToxiGen-RoBERTa [6], a model specifically trained for this dataset. We presented the toxicity scores as averaged probabilities in percentage form.
| Demographic Group | Original Toxicity | T5 | BART | Llama 3.2 -1B | Falcon-7B | GPT-3.5 | GPT-4 | SR (IFT) |
| Women | 92.60 | 25.74 | 24.10 | 4.98 | 13.92 | 3.38 | 1.02 | 3.08 |
| Mental Dis. | 90.45 | 18.27 | 18.29 | 2.25 | 7.28 | 3.65 | 1.12 | 1.12 |
| LGBTQ | 86.58 | 21.89 | 20.01 | 2.34 | 11.13 | 3.24 | 0.67 | 1.78 |
| Black | 90.48 | 26.35 | 26.01 | 2.98 | 12.21 | 4.12 | 1.66 | 1.02 |
| Chinese | 86.52 | 17.68 | 16.74 | 1.99 | 8.23 | 3.25 | 1.46 | 0.75 |
| Asian | 99.19 | 16.77 | 15.10 | 1.73 | 6.02 | 4.10 | 1.23 | 1.29 |
| Native Amer. | 98.27 | 20.96 | 19.35 | 2.12 | 11.62 | 3.81 | 1.47 | 1.83 |
| Middle Est. | 91.54 | 23.47 | 23.45 | 2.18 | 8.34 | 4.06 | 1.01 | 1.83 |
| Muslim | 94.46 | 23.79 | 24.50 | 2.98 | 13.98 | 3.79 | 1.12 | 1.58 |
| Physical Dis. | 82.84 | 18.82 | 17.02 | 1.56 | 6.82 | 3.18 | 0.59 | 1.08 |
| Mexican | 87.48 | 34.27 | 33.56 | 4.01 | 14.21 | 3.80 | 1.22 | 1.20 |
| Jewish | 81.96 | 23.28 | 26.39 | 3.10 | 15.53 | 3.78 | 1.19 | 2.26 |
| Latino | 84.84 | 29.45 | 30.12 | 4.03 | 15.87 | 3.57 | 1.34 | 2.10 |
The results presented in Table 5 indicate that GPT-4, and SR in both variations (prompting and IFT) consistently outperformed other models in generating texts with minimal toxic content across diverse demographic groups. Specifically, GPT-4 recorded the lowest toxicity percentages for groups including women (1.02%), LGBTQ (0.67%), Physical disability (0.59%)and other demographics, as highlighted in Table 5. SR also showed robust performance, especially among groups such as individuals with the Chinese, Mexicans and Black demographics. In contrast, models such as T5 and BART were less effective, exhibiting the highest percentages of unsafe content across most demographic categories.
Main Finding: Closed-source GPT-4 is a strong model in reducing toxicity, while smaller models like T5 and BART struggle in this regard. Instruction fine-tuning models on the top of inherently safe models (e.g. our model SR compared to base model) perform better than simple fine-tuning or prompt-tuning, addressing RQ3.
4.3 Performance evaluation for stereotypes
We assessed SR alongside other leading models using the StereoSet [30] to determine bias across four demographic categories: gender, profession, race, and religion. Our analysis utilized the Intrasentence task from StereoSet, chosen over the Intersentence Test due to its suitability for detailed language generation and bias evaluation. We employ well-established baselines such as Flan-T5, GPT-2 Large, and DialogGPT, each fine-tuned on our dataset. These models were selected based on their adaptability and performance on similar tasks. The goal of this experiment is to see if task-specific instruction fine tuning can reduce biases in text, while retaining language understanding.
| Gender | Profession | |||||
| Model | LMS () | SS | ICAT () | LMS () | SS | ICAT () |
| Flan-T5base FT | 87.84 | 56.70 | 76.07 | 89.01 | 59.64 | 71.85 |
| Flan-T5medium FT | 88.63 | 55.31 | 79.22 | 84.32 | 61.79 | 64.44 |
| Flan-T5large FT | 92.55 | 65.25 | 64.32 | 91.36 | 61.62 | 70.13 |
| GPT2large FT | 80.77 | 70.93 | 46.96 | 79.99 | 64.34 | 57.05 |
| DialoGPTlarge FT | 82.50 | 61.29 | 63.87 | 79.87 | 58.72 | 65.94 |
| Llama3.2 1B (P) | 90.12 | 59.18 | 61.29 | 90.25 | 56.19 | 68.58 |
| SR (IFT) | 91.05 | 54.47 | 76.98 | 91.98 | 62.10 | 74.12 |
| Race | Religion | |||||
| Model | LMS () | SS | ICAT () | LMS () | SS | ICAT () |
| Flan-T5base FT | 86.38 | 68.23 | 54.89 | 83.54 | 69.70 | 50.63 |
| Flan-T5medium FT | 83.47 | 62.52 | 62.57 | 83.54 | 62.12 | 63.29 |
| Flan-T5large FT | 91.48 | 62.16 | 69.23 | 96.20 | 80.26 | 37.98 |
| GPT2large FT | 69.43 | 68.35 | 43.95 | 66.09 | 75. | 31.76 |
| DialoGPTlarge FT | 83.44 | 60.51 | 65.90 | 84.91 | 67.23 | 55.65 |
| Llama3.2 1B (P) | 91.28 | 66.78 | 60.65 | 91.23 | 62.92 | 69.48 |
| SR (IFT) | 93.48 | 52.76 | 74.13 | 93.99 | 62.23 | 77.24 |
The results presented in Table 6 demonstrate the superior performance of SR in reducing stereotypical biases across multiple dimensions: Gender, Profession, Race, and Religion. In particular, SR achieves an impressive balance between reducing biases and preserving language modeling capabilities, which is evidenced by higher ICAT scores across various demographics. SR exhibits balanced performance in the Gender and Profession categories, with high LLMS of 91.05 and 91.98 percentages respectively, while maintaining low SS (close to 50) and high ICAT scores (above 50). In the Race and Religion categories, SR not only achieves high LMS scores (93.48 for Race and 93.99 for Religion) but also records the lowest SS (closer to 50).
Main Finding: SR is an instruction fine-tuned model that reduces stereotype biases while preserving the knowledge and language understanding of the model, addressing the RQ1.
4.4 Performance evaluation on text style factors
Text style factors refer to the distinctive elements that influence the presentation and perception of written text, including vocabulary, sentence structure, tone, voice, and formatting [51]. These factors define the unique character and readability of the text, contributing to how effectively it communicates with specific audiences.
4.4.1 One-sample t-Test for safety measure
To evaluate the effectiveness of SR, we conducted a controlled experiment utilizing the ParlAI style classifier [51] to analyze the stylistic attributes of texts before and after safety interventions (using our instruction fine-tuning method to debias the texts). This experiment aimed to determine whether these safety interventions led to significant changes in linguistic style and CLEN scores. For instance, a higher CLEN value associated with a positive trait like ‘scholarly’ indicates a more consistently positive style. We want to see after debiasing, whether we got higher CLEN with many positive traits.
The experiment was performed using a one-sample t-test, comparing the mean style scores of 16,602 samples from our training set against a hypothesized neutral value, which indicates no unsafe generations. The null hypothesis () assumes that there is no significant change in style due to the our safety measures, while the alternative hypothesis () posits a discernible shift towards safer expressions.
The results, highlighted in Figure 3, demonstrated a statistically significant change in the linguistic style post-intervention, with a p-value of less than 0.00001, leading to the rejection of . The significant t-statistic of 28.17 further confirmed the effectiveness of the safety measures, showing a pronounced improvement in the model output. This result indicates that our approach successfully mitigates bias by removing negative styles, and keep the safety of language generated by SR. It shows a substantial shift in stylistic features towards more safe and inclusive traits.
4.4.2 Stylistic variations post-safety intervention
We also show the effectiveness of safety interventions on SR through style classification on original unsafe text and then benign (safe) generations through our model. Figures 4 show a significant reduction in negative traits and an enhancement of positive attributes in style post-intervention. Figure 4 illustrates contrasting collections of personality traits, each with a different focal point and emotional tone. The first network, shown in Figure 4(a), emphasizes more negative or challenging traits such as “Neurotic",‘ ‘Hostile". and “Cruel". In contrast, the second diagram, shown in Figure 4(b), highlights positive and socially admirable qualities like “Caring", “Compassionate" , and “Honest".
Main Finding: Instruction fine-tuning, as in the case of SR, tends to reduce negative styles in the content, and more positive traits in the language generations, answering the RQ2 and also RQ1.
4.5 Evaluating instruction fine-tuning, prefix-tuning, and prompt-based approaches on bias reduction and knowledge retention
The goal of this experiment is to evaluate whether instruction fine-tuning enhances model performance and bias mitigation more effectively than zero-shot and few-shot prompting methods, to answer our RQ2. We used the Llama3.2-1B instruct model as the baseline for its capabilities and safety features. The settings for this experiment are:
Zero-Shot Prompting: The model is used in a zero-shot setting, where it responds to prompts designed to test bias without any prior specific training on the examples.
Few-Shot Prompting: The model is tested with a few examples before being prompted to respond, allowing it to adapt its responses based on the limited provided context.
Prefix-Tuning: In this approach, a small, task-specific prefix is added to the model’s input. Unlike full fine-tuning, where the entire model’s parameters are adjusted, prefix-tuning modifies only a small set of parameters associated with the prefix, leaving the core model unchanged.
Instruction Fine-Tuning: The model is fine-tuned using our custom dataset, resulting in SR, which includes a balanced mix of biased and unbiased text examples. The fine-tuning process involves adjusting the model’s parameters to better recognize and correct unsafe (biased) content.
To evaluate, we utilized both our in-house test set and the ToxiGen database. For scoring, we employed the LLM-based moderation API provided by OpenAI. Additionally, we calculated the Knowledge Retention metric to assess whether the LLM retains factual information from the input in its generated output.
| Text | Our Test Set | Toxigen Test Set | ||
| OpenAI Mod. | Knowledge Retention | OpenAI Mod. | Knowledge Retention | |
| Pre-safety Scores | ||||
| Original Texts | 57.82% | N/A | 69.78% | N/A |
| Post-safety Scores | ||||
| Baseline (zero-shot) | 21.14% | 64.32% | N/A | 70.19% |
| Baseline (2-shots) | 13.05% | 72.25% | 14.88% | 66.93% |
| Baseline (5-shots) | 12.54% | 73.89% | 12.23% | 79.34% |
| Baseline (Prefix-tuning) | 10.02% | 82.10% | 10.12% | 79.10.34% |
| SR | 05.92% | 90.18% | 05.10% | 87.53% |
The analysis in Table 7 explores the comparative performance of baseline model with prompts (zero, 2, and 5 shots) and our instruction fine-tuned SR model.
Pre-Safety Scores: The original sentences show high moderation scores of 57.82% for our test set and 69.78% for the Toxigen test set, indicating a need for intervention.
Post-Safety Intervention: Post-intervention, we observe a desirable decrease in moderation scores and an increase in knowledge retention as we escalate the number of examples in prompts from zero to 2 to 5 shots and prefix-tuning. Our SR model demonstrates a significant reduction in moderation scores to 5.92% on our test set and 5.10% on Toxigen, respectively, while achieving the highest scores in knowledge retention ( about 90% and 87%, respectively). The performance gap between prefix tuning and instruction fine-tuning on SR is minimal across various test sets. This finding indicates that prefix tuning is a viable and efficient alternative for enhancing model safety, especially in environments with limited computational resources.
Main Finding: Instruction fine-tuning LLMs on custom data can effectively reduce bias and toxicity while retaining substantial knowledge, addressing RQ1. Prompts with demonstrations also proves beneficial. More data improves performance; for example, 5-shot prompts performed better than 2-shot and zero-shot, with instruction fine-tuning performing best, addressing RQ2. Between prefix-tuning and instruction finetuning, we observe marginal difference. The choice of method depends on available compute and whether a more efficient model or one with some better performance is desired. Achieving optimal knowledge retention is best achieved through instruction fine-tuning in this experiment.
4.6 Human Evaluation
The goal of this experiment is to see if our task-specific instruction fine tuning can retain language understanding while reducing biases (RQ1).
We assess SR and its variants for their ability to minimize harm, bias, toxicity, and stereotypes, targeting content neutrality and inclusivity through human evaluation. The model variations include:
Safe_PEFT-1_ep, SR fine-tuned for one epoch and PEFT- our default setting.
Safe_PEFT-5_ep, SR extends PEFT to five epochs.
Safe_Dense-IT or prefix-tuning , SR adjusting prompts on base model without fine-tuning (using dense method) .
Safe_Dense-IFT, SR subjected to dense fine-tuning over five epochs .
A team of five evaluators from our team conducts a rapid, blind assessment of 100 examples of the safe text generations, each contain instances of bias, toxicity, negative sentiments and harm. These variants are tested against criteria based on:
Safety: Ensuring outputs are free from bias, toxicity, prejudice and favoritism and are safe text generations.
Language Understanding: Ensuring that outputs maintain the integrity of the original text content without compromising its meaning or accuracy.
Main results of human evaluation are compiled in Table LABEL:tab:human, summarizing the human evaluation outcomes.
| Prompt | Response |
| Prompt with original text | Revise the provided text to ensure it is neutral, inclusive, and respectful, removing any implicit or explicit biases. Focus on using language that does not favor any particular gender, ethnicity, age group, culture, or ability. Replace any stereotypes with factual, unbiased descriptions, and ensure that all references to individuals or groups are made in a way that respects their dignity and diversity. The objective is to present information in a manner that is accessible and respectful to all readers, promoting equality and understanding. The revised text should maintain the original message’s intent but be free from assumptions, generalizations, and any language that might be considered harmful or exclusive. |
| Response | Successful tech entrepreneurs are usually young men who have a natural talent for understanding complex technologies. |
| Model | Safe_PEFT-1_ep (default SR setup) |
| Response | Successful tech entrepreneurs come from diverse backgrounds, encompassing all genders and ages, and succeed through a combination of skills, knowledge, and dedication. |
| Safety & Lang. scores | 5/5 ; 4.99/5 |
| Insight | This response excellently addresses the prompt by removing any specific gender and age assumptions, promoting a more inclusive view of tech entrepreneurship. It emphasizes diversity, skill, knowledge, and dedication as the keys to success, which aligns perfectly with the aim to neutralize biases. The safety and language scores are justifiably high, reflecting the response’s effectiveness in promoting inclusivity while maintaining clarity and coherence. Text length is quite close to actual text. |
| Model | Safe_PEFT-5_ep |
| Response | Successful tech entrepreneurs are often perceived as hombres jóvenes, though this overlooks the diverse range of individuals who excel in the field. While some may possess habilidad natural for understanding complex technologies, many others from different backgrounds also contribute significantly to the tech industry. However, it’s essential to acknowledge that el sesgo aún existe and may influence our perceptions in this context. |
| Safety & Lang. scores | 4/5; 3.5/5 |
| Insight | This model’s response introduces a change by directly translating some terms into Spanish, which interestingly adds a multicultural dimension but may inadvertently narrow the focus to a specific linguistic or cultural group, potentially conflicting with the goal of full inclusivity. The inclusion of Spanish phrases (’hombres jóvenes’, ’habilidad natural’, ’el sesgo aún existe’) is creative but might not fully address the global nature of tech entrepreneurship. |
| Model | Safe_Dense-IFT |
| Response | Successful tech entrepreneurs are often perceived as young men, though this overlooks the diverse range of individuals who excel in the field. While some may possess a natural talent for understanding complex technologies, many others from different backgrounds also contribute significantly to the tech industry. However, it’s important to recognize that biases still exist…. |
| Safety & Lang. scores | 3.4/5; 3.8/5 |
| Insight | Reiterating the theme of diversity in tech entrepreneurship, this response mirrors the initial corrective approach but concludes abruptly, leaving a sentence unfinished. This might suggest a lapse in addressing the complete scope of the prompt. The ellipsis (’exist….’) could indicate an incomplete thought process or a technical error in response generation. The safety and language scores indicate room for improvement, especially in fully addressing bias and inclusivity. |
| Model | Safe_Dense-IT |
| Response | I’m sorry for any unintentional bias in the previous statement..Tech entrepreneurship thrives on diversity, with successful individuals spanning various demographics. It’s not limited to any particular age, gender, or background. These entrepreneurs leverage a wide array of skills and experiences to navigate complex technologies and drive innovation. |
| Safety & Lang. scores | 4.5/5 ; 4.8/5 |
| Insight | While the apology for any unintentional bias is appreciated, it may be perceived as unnecessary in this context. Nevertheless, this response successfully expands the narrative to encompass a broader range of demographics, emphasizing the diversity and complexity of tech entrepreneurship. The scores indicate a commendable performance in promoting safety and understanding. |
The evaluation of SR model variations in Table LABEL:tab:human demonstrates diverse approaches to enhancing text neutrality and inclusivity. The Safe_PEFT-1_ep model performed best in eliminating unsafe generations (through examples emphasizing diversity and skill in tech entrepreneurship), resulting in the highest scores for safety and language understanding. Conversely, the Safe_PEFT-5_ep model introduced other language (e.g., Spanish) phrases and fell slightly short of achieving full inclusivity. The Safe_Dense-IFT model showed diversity in tech but concluded its responses abruptly, which affected its evaluation scores. Meanwhile, the Safe_Dense-IT model, despite an unnecessary apology for bias, effectively broadened the demographic narrative, scoring well in both safety and understanding.
This qualitative assessment on 100 samples is also shown in Figure 5, where we observe that our default SR method with 1 epoch performs best.
Main Finding: Our current model setup with instruction fine-tuning on our dataset for one epoch performed best. We are able to reduce bias while retaining knowledge.
5 Case Study: Debiasing Job Postings
5.1 Purpose and objective
Online job postings are a critical first point of contact between employers and prospective candidates. Biased or exclusionary language can inadvertently discourage certain applicant groups, which undermines diversity and inclusion efforts. This case study explores how SR mitigates biases in publicly sourced job postings while preserving each posting’s domain relevance and accuracy. Our objective with this case study is to (1) detect and remove potentially biased or gender-coded language, (2) retain essential information about job responsibilities, qualifications, and organizational culture (3) to make the approach be adapted to a large corpus of job postings, even without a direct collaboration with a specific company.
5.2 Methodology
Data collection
We curated a dataset by gathering publicly available job postings accessed via online job boards and search engines, such as Google, focusing on a manageable subset of 500 postings to ensure high-quality analysis and annotation. The postings were sourced from various industries, including Technology, Finance, and Healthcare, and represented roles across different levels (Entry-level, Mid-level, and Senior).
Additionally, we developed a supplementary instruct dataset comprising 200 curated examples of biased text snippets (e.g., “He will lead the project…”) paired with neutral rewrites (e.g., “This role leads the project…”). Each example included a rationale to clarify the reasoning behind the transformation, emphasizing bias reduction while preserving semantic integrity. The data collection and curation process followed ethical research practices and platform guidelines, ensuring compliance and responsible use of publicly available information.
Model building
We used Llama-3.2-1B-Instruct888https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct, a model with 1 billion parameters, which includes enhanced safety guardrails. It was pretrained on a multilingual corpus and fine-tuned for general-purpose English tasks.
SR Pipeline:
Phase 1 (Bias Instruction Training Data): We taught the model to rewrite biased statements by feeding it the instruct dataset.The adoption of the Alpaca-style format further enhanced the model’s ability to follow precise instructions.
Phase 2 (Domain Adaptation): We fine-tuned further on the collection of public job postings with biased-debiased version of the data to ensure the model learned the stylistic and semantic norms of typical hiring announcements.To optimize training efficiency while maintaining effectiveness, technique like QLoRA was employed. The autoregressive loss minimizes the discrepancy between generated responses and ground truth (neutral rewrites), enabling the model to align its outputs with the debiasing goals.
Phase 3 (Iterative Feedback): A panel of five expert reviewers (via LabelStudio) evaluated the model’s debiasing suggestions, approving or rejecting changes based on their quality and relevance. This iterative feedback loop refined the model’s ability to handle ambiguous cases and improved its overall performance.
5.3 Evaluation
Quantitative results
The performance of SR was evaluated using three primary metrics: Bias Reduction, Semantic Fidelity, and Perplexity. These metrics capture the model’s effectiveness in debiasing text while preserving fluency and content fidelity.
1.
We measured bias reduction by calculating the average number of flagged biased terms per job posting before and after fine-tuning.
2.
Using sentence-level embedding similarity, we quantified the semantic overlap between the original and revised texts.
3.
Perplexity, which reflects the fluency and coherence of generated text, was also evaluated. Lower perplexity values indicate better model performance.
| Metric | Original (Baseline) | Post- SR (Fine-Tuned) |
| Avg. Biased Terms/Post | 3.5 | 1.4 |
| Content Similarity (Cosine) | - | ~88% |
| Perplexity | 22.0 | 19.3 |
The results summarized in Table 9 demonstrate the effectiveness of SR in achieving bias reduction, semantic preservation, and improved text fluency. The SR achieved a 60% reduction, decreasing biased terms per post from 3.5 to 1.4 in the test set. The model reduced the average number of biased terms per job posting by over 60%, indicating substantial progress in eliminating biased language. Semantic preservation, measured through content similarity (cosine), showed an 88% similarity score between original and revised texts, ensuring that most domain-specific details were retained with minimal distortion. Additionally, perplexity—a metric reflecting text fluency and coherence—improved significantly, decreasing from 22.0 to 19.3, which highlights the model’s ability to produce more readable outputs. Together, these metrics validate SR as a highly effective solution for debiasing tasks in real-world applications.
Qualitative Assessment
The 5 reviewers scored each revised posting on a 1–5 scale, focusing on inclusivity and clarity. Average ratings rose from 3.6 (original text) to 4.2 (revised text). Several annotators noted improved neutrality when references to age or specific cultural traits were removed. Some flagged instances where domain-specific language was mistakenly altered (e.g., acronyms or technical terms not recognized by the model’s lexicon).
Although SR effectively mitigated many biases, a subset of postings contained subtle stereotypes or domain-specific jargon that wasn’t addressed. Examples include phrases referencing “culture fit,” which can carry implicit biases, and unique role-based terminology sometimes confused for biased language (e.g., “unicorn designer”).
Proof of Concept
This case study underscores the potential of an instruction fine-tuned model to reduce biased language in publicly sourced job descriptions. This approach is reproducible and doable even with fewer compute resources, since we introduced the annotation scheme and the parameter efficienct methods to train the model. This makes the approach more accessible to researchers and smaller organizations. The methodology could extend beyond job descriptions to other high-impact areas such as academic admissions, policy drafts, or technical documentation, all without the need for proprietary corporate data.
6 Discussion
6.1 Theoretical and Practical Implications
This study provides several theoretical insights into the role of curated datasets in mitigating biases within LLMs, contributing to foundational research in AI ethics, fairness, and safety. We examine how instruction fine-tuning with targeted datasets can reduce bias. The work deepens our understanding of bias dynamics and enhances the broader discourse on designing responsible and trustworthy AI systems.
From a practical perspective, the findings emphasize the importance of constructing diverse and inclusive datasets that facilitate equitable model performance across varied contexts. This dataset not only helps in fine-tuning LLMs to produce debiased outputs but also addresses critical challenges in AI interpretability and accessibility. Together, these implications highlight actionable strategies for leveraging datasets as tools to build AI systems that are both technically robust and socially responsible.
6.2 Limitations
Like any study, this work has limitations that must be acknowledged:
Coverage and Diversity of Data
The dataset, comprising annotated news and social media articles, spans a variety of topics and media types. However, it is not fully representative or balanced across all countries, regions, or demographics. This uneven coverage could result in gaps, particularly in capturing nuanced demographic contexts and techniques, limiting the dataset’s applicability to global scenarios.
AI Safety and Emerging Biases
The rapid advancement of LLMs necessitates a strong focus on AI safety. While this study addresses numerous known challenges, the continuous evolution of AI technologies introduces new complexities and unforeseen risks. Effectively accounting for these emerging issues remains an ongoing challenge.
Bias and Subjectivity
Bias is an inherent and multifaceted issue in any dataset [52]. Despite implementing detailed annotation guidelines, the subjectivity of annotators and evaluators cannot be entirely eliminated. Systemic biases within the source data may also persist, and efforts to address demographic and contextual biases are unlikely to cover all possible manifestations. This limitation underscores the need for continual refinement and vigilance.
Methodological Constraints
Training and evaluating LLMs demand significant computational resources, posing accessibility challenges for smaller research groups. While optimization methods such as PEFT and QLoRA are used to alleviate resource demands, these approaches require specialized expertise and can introduce additional complexity. Dense fine-tuning, while explored, does not always yield consistent performance improvements. Moreover, reliance on proprietary evaluation platforms, such as OpenAI APIs requiring access keys, restricts flexibility, reproducibility, and transparency in model assessment.
6.3 Future Directions
Addressing the limitations identified in our study, future research should aim to curate more globally representative datasets and enhance AI safety protocols to adapt to evolving challenges. Developing sophisticated bias mitigation strategies is also crucial. Methodological advancements that lower computational demands and simplify the optimization process are needed to enable wider accessibility for diverse research groups. Furthermore, establishing open, flexible evaluation frameworks and engaging in discussions around ethical considerations and potential regulatory frameworks are essential. These efforts will collectively advance the development of safe and responsible LLMs that align well with societal values.
While the proposed method has shown effectiveness in enhancing the model’s safety and knowledge retention, it might be worthwhile to assess its impact on the model’s general capabilities in various NLP tasks. Future work should focus on conducting comparative evaluations using datasets like Massive Multitask Language Understanding (MMLU) [53]. These evaluations will provide insights into whether the safety improvements lead to any significant trade-offs in the model’s general performance.
In future work, we aim to enhance the reliability of our bias assessment by conducting a larger number of trials and employing advanced statistical techniques, such as chi-square tests, to robustly analyze gender representation in model outputs. This will enable a more definitive conclusion about the presence of gender or other biases.
7 Conclusion
In this study, we introduced SR for safe language generations, this approach is trained on our custom dataset of instructions featuring original texts (potentially unsafe) and their benign variations to ensure safe language generation. This model offers reduced inference and deployment costs. It has proven competitive in many benchmarks. We have detailed the methods and techniques to develop our models, emphasizing their adherence to safety and language understanding principles. Committed to transparency and safety, we plan to enhance the model and data in future work.
Acknowledgements Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.
Declarations
Funding: No funding is available.
Conflict of Interest: The authors have no conflict of interest to declare that are relevant to the content of this article.
Ethics Approval and Consent to Participate: Not applicable.
Consent for Publication: Not applicable.
Data Availability: Data is made publicly available in Huggingface.
Code Availability: The code is made publicly available at GitHub.
Author Contribution:
Shaina Raza: Formal analysis, Data curation, Conceptualization, Methodology, Experiments, Visualization, Writing – original draft, Writing – review & editing, Supervision.
Oluwanifemi Bamgbos: Investigation, Methodology.
Shardul Ghuge: Experiments, Validation, Visualization, Review.
Fatemeh Tavakoli: Investigation, Experiments.
Deepak John Reji: Statistical analysis.
Syed Raza Bashir: Validation, Experiments, Writing – review & editing.
Appendices
Appendix A Annotation Guidelines
In the development of this guide, a dedicated group of 20 annotators volunteered their expertise and time to ensure the highest standards of accuracy and sensitivity in identifying unsafe content generation. This diverse team consisted of five experts in fields related to computer science, language, psychology, and ethical computing, each accompanied by four students. An annotation guide is designed for the annotators for identifying and effectively neutralizing instances of toxicity, stereotyping, bias, and harm in textual content. This guidelines covers a diverse array of target groups and individuals, ranging from those in different age groups – children, teenagers, adults, and seniors – to individuals with varying educational backgrounds, geographic locations, and occupations, including healthcare professionals, engineers, teachers, and artists. Our aim with this annotation guideline is to ensure that all individuals, regardless of their background or characteristics, can engage with content that adheres to the principles of fairness, inclusivity, and respect. Below is a list of target groups for which unsafe language generation happens that need to consider when annotating textual content and creating benign alternatives.
Target Group Attributes:
Age Group: Children, Teenagers, Adults, Seniors;
Education Level: High School, College, Postgraduate;
Geographic Location: Urban, Suburban, Rural;
Occupation: Healthcare Professionals, Engineers, Teachers, Artists;
Language: English Speakers, Spanish Speakers, Mandarin Speakers;
Hobbies & Interests: Sports Enthusiasts, Music Lovers, Travel Enthusiasts;
Family Status: Singles, Married, Parents, Divorced;
Technology Adoption: Early Adopters, Technophobes;
Environmental Concerns: Eco-conscious, Sustainability Advocates;
Dietary Preferences: Vegetarians, Vegans, Meat Lovers;
Political Affiliation: Conservative, Progressive;
Religion: Buddhism, Catholic, Christian, Islam;
Race & Ethnicity & Nationality: Asian, Black, Chinese, Indian, Korean-Chinese, Southeast Asian, White;
Gender & Sexual Orientation: LGBTQ+, Men, Women;
Miscellaneous: Ageism, Disabled, Diseased, Feminist, Physical Appearance, Socio-economic Status.
Following are the examples for annotation guidelines.
Bias manifests as the display of favoritism or prejudice towards certain individuals, groups, or ideas, often leading to unfair treatment or judgment. It involves a preference for one perspective or the unfair exclusion of others, thereby distorting objectivity and creating an unequal or discriminatory impact.
An example of bias is : Clearly, anyone with a real education would naturally steer clear of such backward political views. It is obvious that educated individuals know better. This statement is biased, saying that educated people uniformly reject certain political views. Annotators should identify the assumption of uniformity in thought among educated people, highlight the biased words, as: real education, backward political views, educated individuals know better, recommending a more inclusive statement like, “Individuals of all educational backgrounds can hold a variety of political views, reflecting the diversity of thought and opinion".
Toxicity encompasses language that not only exhibits harm, aggression, or disrespect but also creates an atmosphere of hostility, offense, or harm towards individuals or groups. It includes expressions that go beyond civil discourse and can lead to emotional or psychological harm by targeting someone’s identity, intelligence, or worth.
Consider an example: It is baffling how anyone with half a brain could possibly agree with such a ludicrous stance. Only an idiot would fail to see the flaws". Here, the language is blatantly toxic, attacking the intelligence of those with differing opinions. This sentence should be labeled as ‘toxic’. When annotating, phrases like anyone with half a brain , Only an idiot should also be marked as toxic. A benign revision might be, “This stance may not be universally accepted, as it has several notable flaws that could be overlooked”.
Stereotyping occurs when sweeping generalizations are made about a particular group of people, often based on limited or inaccurate information. These generalizations tend to be overly simplistic and fail to account for the diversity and individuality within the group, resulting in a misleading or unfair portrayal. Specifically, we focus on the negative aspects of sentiment associated with such stereotypes.
An example might be, Every tech enthusiast I have met so far is socially awkward, buried in their gadgets, and incapable of holding a meaningful conversation about anything else. This stereotype unjustly labels a diverse group. Annotators should highlight the entire stereotypical portrayal, highlighting phrases socially awkward,buried in their gadgets,incapable of holding as conveying stereotypes, suggesting a benign variation such as, “Many tech enthusiasts are deeply passionate about technology, yet also possess a wide range of interests and social skills”.
Harm encompasses content that has the potential to inflict distress, emotional pain, or harm to individuals or communities. It may also involve the explicit or implicit advocacy for violence, whether physical, emotional, or psychological, against individuals, groups, or entities. This category includes expressions that can have severe negative consequences on individuals’ well-being or safety.
An example: In times like these, violent actions are not only understandable but fully justified. Those who stand in the way of progress deserve to face the consequences. This example advocates violence as a response to conflict. Annotators need to recognize the advocacy for violence, highlighting words like violent actions are… fully justified, suggesting a benign variation that promotes non-violent advocacy: “There are various ways to confront these issues, including peaceful protest and dialogue, which can lead to constructive outcomes without resorting to violence”.
Annotation Process & Ethical Considerations
Annotators should: Carefully read the text to identify any instances of toxicity, stereotyping, bias, or harm. Use the guidelines to determine the best way to annotate these instances. Suggest benign variations that maintain the original message’s intent without the biased content. Annotators must remain neutral and respectful, considering the impact of words on diverse audiences. Ethical annotation respects cultural differences and promotes inclusivity.
Training, Resources, & Feedback Mechanism Ongoing education is crucial. Annotators are encouraged to engage with training materials and participate in workshops to refine their skills. Additionally, annotators should provide feedback on the guidelines, share insights from challenging texts, and suggest improvements to ensure the guidelines evolve to meet emerging needs.
Inter-Annotator Agreements
Fleiss’ Kappa is a statistical measure used to assess the agreement among multiple raters or judges when categorizing items into multiple classes. It quantifies the level of agreement beyond what would be expected by chance alone. The score interpretation is as:
less than 0.21 Poor
0.21-0.40: Fair agreement
0.41-0.60: Moderate agreement
0.61-0.80: Substantial agreement
0.81-1.0: Almost perfect agreement
| Category | Fleiss Kappa Score | Interpretation (Agreement) |
| Toxicity | 0.75 | Substantial |
| Harm | 0.68 | Substantial |
| Bias | 0.62 | Substantial |
| (-ve) Sentiment | 0.70 | Substantial |
| Benign variation | 0.78 | Substantial |
The Fleiss Kappa scores, a measure of inter-annotator agreement, were thoroughly calculated for each category. For toxicity, we observed a Fleiss’ Kappa score of 0.75, indicating strong agreement. The harm category showed a substantial consensus with a score of 0.72. Bias identification and removal also achieved a high degree of agreement, with a score of 0.78. Negative sentiment analysis resulted in a good agreement score of 0.70. The benign generation score is calculated to be 0.78.
An annotated example is given below:
- 1001: Millennials in the Workplace
Original: Millennials are too sensitive and don’t understand real work.
Benign: There is a perception that millennials have different work ethics, which may be misunderstood.
Details: Bias - Yes, Toxicity - Yes, Sentiment - Negative, Harm - Medium, Target - Millennials
Biased Words: "too sensitive", "don’t understand real work" - 1002: Women in STEM
Original: Women are not as good in STEM fields as men.
Benign: People of all genders can excel in STEM fields.
Details: Bias - Yes, Toxicity - No, Sentiment - Negative, Harm - Medium, Target - Women
Biased Words: "not as good" - 1003: Language Diversity
Original: I hate it when people speak Spanish in America. Speak English!
Benign: It’s important to respect linguistic diversity in America.
Details: Bias - Yes, Toxicity - Yes, Sentiment - Negative, Harm - High, Target - Non-English Speakers,
Biased Words: "hate", "Speak English!"
Appendix B Analysis of the CMD - Dataset used in this Study
B.1 Data Schema
B.2 Data Analysis
| Statistic | char_length | word_length |
| Count | 16614.000 | 16614.000 |
| Mean | 370.336 | 68.277 |
| Std | 451.732 | 80.136 |
| Min | 7.000 | 1.000 |
| 25% | 110.000 | 22.000 |
| 50% | 238.000 | 44.000 |
| 75% | 491.000 | 90.000 |
| Max | 5000.000 | 1108.000 |
Table 11 shows the text length statistics indicate moderate average lengths but with considerable variability, suggesting the dataset contains a diverse array of text sizes, which is advantageous for creating versatile NLP models. The labels distribution is in Appendix12.
| Label | Category (Count) |
| BIAS | No (14227), Yes (5772) |
| TOXICITY | No (12040), Mild (5293), High (2666) |
| SENTIMENT | Negative (9028), Neutral (8370), Positive (2601) |
| HARM | Low (14151), Medium (3932), High (1915) |
Figure 6 Flesch Reading Ease scores imply that positive sentiments are found in more straightforward texts, while negative and neutral sentiments appear in more complex ones, potentially affecting the ease of automated readability assessment.
| Statistic | Value |
| Count | 8307.000000 |
| Mean | 0.087286 |
| Std | 0.148290 |
| Min | 0.000000 |
| 25% | 0.028708 |
| 50% | 0.052493 |
| 75% | 0.084249 |
| Max | 1.000000 |
Table 13 demonstrates the benign analysis of text alterations post-de-biasing, revealing a notable variation in the extent of change. This highlights the complexity of the debiasing process and the need for careful consideration in maintaining the original text content.
Appendix C QLoRA Details
QLoRA (Quantization with Low-Rank Adapters) [27] is an efficient fine-tuning technique used in the context of LLaMA-2. LLaMA-2 is a powerful pre-trained model that can be fine-tuned for various natural language understanding tasks. Fine-tuning involves adapting a pre-trained model to a specific downstream task by training it on task-specific data. It helps the model specialize and perform well on specific tasks. QLoRA is the key technique used during fine-tuning. QLoRA quantizes the pre-trained LLaMA-2 model to use fewer bits for its weights, specifically reducing the precision of the model’s weights to 4 bits. After quantization, QLoRA attaches small “Low-Rank Adapters” to the model, which are fine-tuned on the task-specific data.
Greedy Decoding We use the greedy decoding strategy for generating word-tokens in the inference phase.
Packages We utilize models through HuggingFace Transformers [54], and utilize the Trainer from this package to train our models. In addition, we use integrated bitsandbytes for Parameter Efficient Finetuning.
Appendix D Carbon Footprint Estimation
Quantifying the carbon footprint [39] of machine learning (ML) model training involves calculating the energy consumption during the training process and converting this figure into carbon dioxide equivalent (CO2e) emissions. This calculation is informed by the power usage of the computing hardware (GPUs and CPUs) and the carbon intensity of the electrical supply, which can vary significantly by geographic location and energy source. For our estimations, we adopt a global average carbon intensity of about 0.4 kgCO2e per kilowatt-hour (kWh) as a standard metric.
In evaluating the carbon footprint for training sessions of the SR model, we factored in the energy demands of the hardware components, the length of each training session, and the prevailing carbon intensity of the electricity. Specifically, the PEFT training setup, comprising one A40 GPU and four CPUs for a 50-minute stint, consumed 0.53 kWh of energy, translating to a carbon emission of 0.21 kgCO2e under the assumed carbon intensity rate.
Energy Consumption Metrics:
The process for calculating energy consumption in kWh encompasses the following steps:
- Total Power Consumption: Summation of the GPUs’ and CPUs’ power usage in Watts.
- Session Duration in Hours: Translation of the training period from minutes to hours.
- Calculated Energy Use (kWh): Multiplication of the total power usage (in kW) by the session duration in hours, yielding the total energy consumption.
Carbon Emission Estimation:
The carbon footprint, expressed in kgCO2e, is ascertained by multiplying the total energy consumed (kWh) by the carbon intensity coefficient (0.4 kgCO2e/kWh). The PEFT model training, lasting 50 minutes, necessitated 0.53 kWh of energy, presenting in a carbon emission footprint of 0.21 kgCO2e.
Appendix E Evaluation
E.1 Classifiers
OpenAI Moderation API [45]employs a GPT-based, multi-label classifier specifically fine-tuned to evaluate whether text breaches one of eleven content safety categories. These categories encompass hate, harassment, self-harm, sexual content involving minors, violence, and more. For each category, the endpoint delivers a probability score and a binary label, alongside an overall binary label indicating the content’s safety status.
The Perspective API [44]serves as a tool for online platforms and publishers to identify and remove harmful or offensive content, particularly within comments and discussions. Leveraging machine learning models, it analyzes text to provide probability scores indicating the likelihood of it being perceived as harmful. The risk categories assessed by Perspective API include toxicity, severe toxicity, identity attack, insult, profanity, and threat.
Style Classifier [51]: The style classifier rom ParlAI identifies specific attributes or patterns in text, such as sentiment, formality, or writing style, it is a pre-trained model. In the context of CLEN (Content-Length Entropy Normalization), is a metric that measures the diversity of content length in generated text. Higher CLEN values indicate more varied sentence lengths, which can impact readability and style coherence.
E.2 Safety Prompts
Appendix F Glossary of Terms
Generative AI: AI systems that generate new content or data based on learned patterns.
LLMs: Large Language Models, which are AI models trained on vast amounts of text data to understand and produce human language.
LM: Language Model, a model that predicts the likelihood of a sequence of words.
Misinformation: False or misleading information spread without malicious intent.
Bias: Prejudice in favor of or against one thing, person, or group compared with another, often in a way considered to be unfair.
LLM Alignment: The process of ensuring LLMs’ outputs align with human values and intentions.
Toxicity: The quality of language outputs that could be considered offensive, harmful, or inappropriate.
Stereotype: Oversimplified generalizations about a group that may lead to biased judgments.
Safety in AI or LLMs: Measures and practices to ensure AI systems operate without causing harm or undesired effects.
Guardrails: Predefined rules or limits that guide the safe operation of AI models.
Red-teaming: A practice of challenging a system, model, or organization by simulating potential adversaries.
Data Augmentation: Techniques to increase the diversity of data available for training models without actually collecting new data.
Reinforcement Learning from Human Feedback (RLHF): A technique where models learn from human feedback to improve their performance or alignment with human values.
Instruction Tuning: Fine-tuning AI models on a diverse set of instructions to improve their ability to follow specific commands or intents.
Instruction Fine Tuning: A more focused form of instruction tuning to refine models’ responses to particular instructions.
Safety Context Distillation: The process of condensing and incorporating safety-related information into AI models to guide their outputs.
Prompt Injection: Techniques to influence or control AI model outputs by inserting specific instructions or data into the input.
Adversarial Demonstrations: Examples designed to challenge or trick AI models into making errors, used to improve their robustness.
Safety Intervention: Actions taken to modify an AI system’s design, training, or operation to increase its safety.
Fairness: The principle of making unbiased decisions or predictions, ensuring equitable treatment for all groups.
Quantization: Reducing the precision of the model’s parameters to decrease its size and increase inference speed.
QLORA: A method for quantizing transformers in NLP to reduce model size while retaining performance.
PEFT: Parameter-Efficient Fine-Tuning, techniques that allow for significant model updates without altering the entire model architecture.
Instruction Dataset: A dataset specifically designed for training AI models to follow instructions or perform tasks based on textual commands.
Parameter-Efficient Fine-Tuning (PEFT): PEFT is a technique that allows fine-tuning of large language models like LLama 2 without updating all of the model’s parameters. Instead, it focuses on a small subset of the model’s parameters, making the fine-tuning process more efficient and less resource-intensive. It is efficient and faster fine-tuning
Dense Fine-Tuning (Alternative to PEFT): it is a more traditional approach where all model parameters are updated during fine-tuning. It allows comprehensive adaptation to the new task, but may be slower compared to PEFT.